A 78-nFnF parallel-plate capacitor has plates that sandwich a dielectric layer that is 75 μmμm thick.

Calculate the thickness of the dielectric layer that would reduce the capacitance to 63 nFnF.

Express your answer to two significant figures and include appropriate units.
Answer #1

given

intial capacitance ci = 78*10^-9 F

thickness ti = 75*10^-6 m

from the relation

capacitance c = A/t

here area and medium same

ci*ti = cf*tf

tf = 78*10^-9*75*10^-6/63*10^-9

tf = 92.86*10^-6 m